OpenStudy (anonymous):

What I got when I tried to solve this problem: a[n]= a [1] +(n-1) *d a[n]=14 [1] +( -1) *4 a[n]= 14 + ____*4 What is the simple formula for the nth term of the following arithmetic sequence? 14, 18, 22, 26, 30, … The correct answer is: 4n +10 Please explain how you figure that out

OpenStudy (anonymous):

a1=14 \[a _{n}=a _{1}+\left( n-1 \right)d\] \[a _{n}=14+\left( n-1 \right)4=14+4n-4=4n+10\]

OpenStudy (anonymous):

d=18-14=4

OpenStudy (anonymous):

yep @surjithayer